OpenStudy (anonymous):

x^(2/3) + y^ (2/3) = 4 how would you find an equation for the tangent line at the point (-3sqrt3, 1) using implicit differentiation?

OpenStudy (anonymous):

start with \[\frac{2}{3}x^{-\frac{1}{3}}+\frac{2}{3}y^{-\frac{1}{3}}y'=0\] then replace x and y by the numbers you are given and solve for \[y'\]

OpenStudy (anonymous):

you get a decimal for the slope if you do that though :/

OpenStudy (anonymous):

i agree, you can find a decimal approximation. that is not surprising is it?

OpenStudy (anonymous):

the answer isn't suppose to contain a decimal according to our assignment